3 A recent survey by a local municipality established that daily water usage by its households is normally distributed with a mean of 220 liters and a standard deviation of 45 liters.



(i) What percentage of households is likely to use more than 300 liters of water per day?


(ii) What is the probability of finding a household that uses less than 100 liters of water per day?


(iii) What percentage of households is likely to use between 300 to 350 liters of water per day?



1
Expert's answer
2022-06-15T16:58:39-0400

(i)


"P(X>300)=1-P(Z\le\dfrac{300-220}{45})"

"\approx1-P(Z\le1.7778)\approx0.0377"

"3.77\%"


(ii)


"P(X<100)=P(Z<\dfrac{100-220}{45})"

"\approx P(Z<-2.6667)\approx0.0038"

"0.38\%"


(iii)


"P(300<X<350)=P(Z<\dfrac{350-220}{45})"

"-P(Z\le\dfrac{300-220}{45})\approx P(Z<2.8889)"


"-P(Z\le1.7778)\approx0.0358"

"3.58\%"